The length of a rectangle is 13 inches greater than its width. Its perimeter is 8 ft. Find its dimensions
Answer by addingup(3677)

L = W+13
2L + 2W = 96
substitute for L:
2(W+13) + 2W = 96
2W + 26 + 2W = 96
4W = 70
W = 17.5
-----------------------
Check:
L = 17.5+13 = 30.5
2(30.5) + 2(17.5) = 96 Correct